# Sommerfeld Enhancement for Vector Dark Matter in the Fundamental Representation of $SU(2)_{L}$

### v3

@@ Abstract
In this work, we study an extension of the Standard Model which include a new massive vector field in the fundamental representation of $SU(2)_{L}$. The neutral component of the new vector field is a natural dark matter candidate. We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events. We found that the model exhibits a resonant mass part around the range of $2-10$ TeV. This resonance is influenced by the choice of portal Higgs coupling parameter. We show that these prediction can be partially tested by CTA in the near future.
@@ 1 Introduction
An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet. This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [3] and have been studied in detail for scalar field [4] and fermionic fields [5].
In the case of vector DM (VDM), an effective model including a VDM candidate originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [6] while an ultraviolet complete model was developed in [7]. On the other hand, models including a vector DM candidate coming from an electroweak doublet were studied in  [8, 9]. These vector models are consistent with experimental data, if the new fields has mass in the range of some few TeV. In this mass range, direct detection [10] becomes difficult and the discovery at colliders seems to be challenging even for future colliders [6].
In this circumstances, indirect search can be a privileged alternative to discover or rule out these models. However, it is well known that to obtain an accurate value of the DM annihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
The SE effect was computed for the cases of scalar DM [4, 11, 12, 13, 14], fermion DM field [5, 15] in the doublet and triplet representation of $SU(2)_{L}$ and vector DM in the adjoint representation  [7]. In this work, we focus on computing the SE for the annihilation of VDM in the fundamental representation of $SU(2)_{L}$ and the photon flux (relevant for Indirect Detection  [10, 16, 17, 18]) produced by pair annihilation of the VDM. As we will see in the following section, a unique feature of our case is the presence of non-minimal interactions between the massive vector doublet and the standard gauge bosons.
Our work is organized as follows. In section 2, we describe the main aspects of the model. Section 3 is devoted to the computation of the electroweak, annihilation matrices and annihilation cross section with SE. In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results. Finally, in section 5 we state our conclusions.
@@ 2 The Model
Notice that the last two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons. The appearance of this non-minimal dimension 4 terms is a special feature of having introduced a vector doublet. Similar non-minimal interactions in the scalar and fermion cases happen to be higher dimensional. In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal. For the purpose of comparing, we work with two values of $\kappa=\kappa_{1}=\kappa_{2}$: $\kappa=0$ and $\kappa=1$. The last value coincides with the choice made in [8].
Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet. Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters. However, there is no reason to think that an important hierarchy arise among these couplings. Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$. In this way, $\lambda$ will be the only constant regulating the Higgs portal. After the electroweak symmetry breaking, the physical mass of $V_{\mu}$ ($M_{V}$) receives contributions from the Proca term ($+m_{V}^{2}V_{\nu}^{\dagger}V^{\nu}$) but also from the interactions with the Higgs field.
@@
Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs in the Lagrangian. That means that the Lagrangian is invariant by a $Z_{2}$ transformation under which the new vector is odd and all the standard fields are even. Consequently, this accidental $Z_{2}$ symmetry makes the lighter component of $V_{\mu}$ stable. This stable component happens to be the neutral one ($V_{\mu}^{0}$)
@@ 3.1 Basic Formalism
The so called “Sommerfeld enhancement” is a non-perturbative effect due to the resumation of initial states interactions. This effect may produce sensible consequences for the annihilation cross section in the non-relativistic regime [5, 19]. In the next few paragraphs we will review the formalism used to compute the Sommerfeld enhancement effect [20] and the we will apply it to our model.
@@ 3.1.1 The Two-Particle State and Schrödinger Equation
where index $a$ represents the initial state of the process and $b$ the final state and $r$ is the relative position. The wave function $g_{ab}$ describes the two particles in a doublet of $SU(2)_{L}$. As we have seen, the potential includes interactions with gauge and Higgs bosons, corresponding to Coulomb (long-range) and Yukawa (short-range) potentials. The kinetic energy of these particles is $E\approx\frac{M_{V}v^{2}}{4}$, where $v$ is the relative velocity between them. The $\Gamma$ term is the imaginary potential and corresponds to an absorptive part, arising in the low-energy limit of quartic interactions. Finally, $N$ is the number of spin states, given by $N=(2l+1)^{2}$, which in our case is 9.
@@ 3.4 Numerical Results
Our results are shown in figures 2 and 3 . The first one shows our predictions without taking into account the non-minimal interactions ($\kappa=0$) while in the second figure we used $\kappa=1$. In both cases, there are several noteworthy characteristics of the Sommerfeld enhancement that merit discussion. Firstly, as usual, the SE is a velocity-dependent phenomenon. For instance, no structure is observed for a relative velocity of $v=0.3$ (dashed lines in figures 2 and 3), which is representative of the early Universe, when the relic density was formed. In contrast, for smaller velocities, like the one expected in the galaxy ($v=10^{-3}$), significant oscillations and resonances appear. We have observed that the small oscillations dependent on interactions mediated by the Higgs boson. Indeed, we see that they are attenuated for large values of $\lambda$ parameter Interestingly, the presence of the non-minimal interactions introduce an additional resonance at a lower mass. However, the $\kappa$-terms don’t modify significantly the size of the oscillations nor the amplitude of the resonances. For masses that exhibit such resonances (roughly around $1.9$ and $6$ TeV), the Sommerfeld factor experiences a significant boost of about 5 order of magnitude, for all the values of portal coupling constant $\lambda$.
@@ 4 Photons Flux from Annihilation neutral Dark Matter particles
Using the expressions above, we calculate the photon flux resulting from the annihilation cross section (8), which serves as a method for the indirect detection of DM. This entails the integration of equation (29). Figures 4 and 5 depict the photon flux from VDM annihilation per second for $\kappa=0$ and $\kappa=1$ respectively. In all the cases we have used $\lambda=0.5$. We present the signals of photon flux resulting from the annihilation of VDM particles from the center of the galaxy for different final states ($\gamma\gamma$, $W^{+}W^{-}$, $ZZ$, and $hh$), with a solid angle of $1^{\circ}$, compared with the canonical cross-section with a mass of 100 GeV. Additionally, we show the total flux predicted at different values of the DM mass. We selected the masses to coincide with peaks and local minimums in the SE factor. We compare our results with Fermi-LAT sesitivity and the projected CTA sensitivity [22].
Figure 4: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$. This photon flux is measured at the center of the galaxy, with $\kappa=0$. We show differential sensitivity of CTA [22] and Fermi-Lat [23] in the galactic center.
Figure 5: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, with a Line of sight (los) angle in longitude (llos) of $1^{\circ}$, and latitude (blos) of $0^{\circ}$. This photon flux is measured at the center of the galaxy, with $\kappa=1$. We show differential sensitivity of CTA [22] and Fermi-Lat [23] in the galactic center.
@@ 5 Summary
In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamental representation of $SU(2)_{L}$. We found that the annihilation cross-section predicted by our model is generally bigger than the one predicted in the cases of a spin $1/2$ doublet and a spin $1$ triplet. Additionaly, the presence of non-minimal interactions of the doublet with the standard gauge model produces a richer resonance structure. In fact, in the absence of the non-minimal interaction, only one big resonance in $\sigma v$ is observed for $M_{V}\approx 6$ TeV. However, when the non-minimal interactions are present, a second peak appears at $M_{V}\approx 2$ TeV, in the low velocities regime.
As a product of our work, we used our results on $\sigma v$ to predict the photon flux produced by DM annihilation at the center of the galaxy.

### v4

@@ Abstract
In this work, we study an extension of the Standard Model which includes a new massive vector field in the fundamental representation of $SU(2)_{L}$. The neutral component of the new vector field is a natural dark matter candidate. We compute the Annihilation Cross Section with Sommerfeld Enhancement and the Photon Flux arising from Dark Matter annihilation events. We found that the model exhibits resonances for masses in the range of $2-10$ TeV. This resonance is influenced by the choice of portal Higgs coupling parameter. We show that these prediction can be partially tested by CTA in the near future.
@@ 1 Introduction
An intriguing realization of the WIMP idea is to postulate the DM as the neutral component of an electroweak multiplet. This paradigm, usually called “Minimal Dark Matter”, was proposed long ago [3] and has been studied in detail for scalar fields [4] and fermionic fields [5].
In the case of vector DM (VDM), an effective model that included a candidate VDM originated from a vector field in the adjoint representation of $SU(2)_{L}$ was first proposed in [6], while an complete ultraviolet model was developed in [7]. On the other hand, models including a vector DM candidate coming from an electroweak doublet were studied in  [8, 9]. These vector models are consistent with experimental data, if the new fields have a mass in the range of some few TeV. In this mass range, direct detection [10] becomes difficult, and discovery in colliders seems to be challenging even for future colliders [6].
In these circumstances, indirect search can be a preferred alternative to discover or rule out these models. However, it is well known that to obtain an accurate value of the DM annihilation cross section in the non-relativistic regime, it is necessary to take into account non-perturbative effects like the so called Sommerfeld Enhancement (SE).
The consequences of SE have been extensively studied in various scenarios, including scalar dark matter (DM)[4, 11, 12, 13, 14], fermionic DM fields[5, 15] in the doublet and triplet representations of $SU(2)_{L}$, and vector DM in the adjoint representation [7]. In this work, we focus on computing the SE for the annihilation of vector DM (VDM) in the fundamental representation of $SU(2)_{L}$ and analyzing the resulting photon flux, which is relevant for indirect detection [10, 16, 17, 18].
A distinctive feature of our model is the presence of non-minimal interactions between the massive vector doublet and the standard gauge bosons, which introduce novel dynamics in the annihilation process. Moreover, because the DM particles in our scenario are significantly heavier than the $W$ and $Z$ bosons, the SE effect becomes particularly important and impactful.
Our work is organized as follows. In section 2, we describe the main aspects of the model. Section 3 is devoted to the computation of the electroweak, annihilation matrices, and annihilation cross section with SE. In section 4 we apply our results to compute the photon flux from DM annihilation at the center of the Milky Way and describe our results. Finally, in section 5 we state our conclusions.
@@ 2 The Model
Notice that the last two terms of Lagrangian (2) describe non-minimal interactions between the new vector field and the gauge bosons. The appearance of this non-minimal dimension 4 terms is a special feature of having introduced a vector doublet. Similar non-minimal interactions in the scalar and fermion cases happen to be higher dimensional. In order to avoid a direct photon interaction with $V_{\mu}^{0}$, the otherwise arbitrary constants $\kappa_{1}$ and $\kappa_{2}$ have to be equal (where $\kappa_{1}$ and $\kappa_{2}$ are parameters analogous to couplings in the context of leptoquark vector models). For the purpose of comparing, we work with two values of $\kappa=\kappa_{1}=\kappa_{2}$: $\kappa=0$ and $\kappa=1$. The last value coincides with the choice made in [8]. This model is a simple extension that can also explain the mass of neutrinos, which are not included in the SM.
Another important sector of the Lagrangian is the part of the potential which describes the interaction between the Higgs field and our massive vector doublet. Of course, the coupling constant that regulates the interactions between the Higgs field and the new vector field are free parameters. However, there is no reason to think that an important hierarchy arises among these couplings. Consequently, and for the sake of simplicity, we will take the assumption that all of them are of the same order $\lambda_{2}\approx\lambda_{3}\approx\lambda_{4}\approx\lambda$. In this way, $\lambda$ will be the only constant regulating the Higgs portal. After the electroweak symmetry breaking, the physical mass of $V_{\mu}$ ($M_{V}$) receives contributions from the Proca term ($+m_{V}^{2}V_{\nu}^{\dagger}V^{\nu}$) but also from the interactions with the Higgs field.
@@
Due to a fortunate conspiracy between the Lorentz and gauge structures of the model, the new vector field can only appear in pairs in the Lagrangian. That means that the Lagrangian is invariant by a $Z_{2}$ transformation under which the new vector is odd and all the standard fields are even. Consequently, this accidental $Z_{2}$ symmetry makes the lighter component of $V_{\mu}$ stable. This stable component happens to be the neutral one ($V_{\mu}^{0}$).
@@ 3.1 Basic Formalism
We now compute the indirect signals of $\gamma$-rays produced by dark matter (DM) pair annihilation in specific astrophysical environments, with a particular focus on the Galactic Center of the Milky Way. Our analysis includes the non-relativistic effect known as the Sommerfeld enhancement, which accounts for the resummation of interactions between initial states. This effect can significantly amplify the annihilation cross-section, making it a crucial factor in indirect DM searches [5, 19].
Recent observational data from Fermi LAT and CTA have provided stringent constraints on $\gamma$-ray signals associated with DM annihilation in the Galactic Center, offering invaluable insights into DM properties and annihilation cross-sections [20]. These observations underscore the importance of incorporating the Sommerfeld enhancement in theoretical models to accurately predict $\gamma$-ray fluxes.
In the following sections, we review the formalism employed to calculate the Sommerfeld enhancement effect [21]. We then apply this framework to our model, evaluating its implications for indirect detection strategies using $\gamma$-ray observatories such as Fermi LAT and CTA.
@@ 3.1.1 The Two-Particle State and Schrödinger Equation
The index $a$ represents the initial state of the process, while $b$ corresponds to the final state, and $r$ denotes the relative position. The wave function $g_{ab}(r)$ describes the two particles within a doublet of $SU(2)_{L}$. As previously discussed, the potential $V(r)$ originates from trilinear interactions between the new vector bosons and the electroweak bosons, including the Higgs boson. This can be expressed as $V(r)=V_{gauge}(r)+V_{Higgs}(r)$, incorporating Coulomb (long-range) and Yukawa (short-range) interactions.
The kinetic energy of these particles is approximately $E\approx\frac{M_{V}v^{2}}{2}$, where vv is their relative velocity. The term $\Gamma$ represents the imaginary part of the potential and corresponds to an absorptive component arising from quartic interactions in the low-energy limit. Lastly, $N$ denotes the number of spin states, given by $N=(2l+1)$, which in our specific case evaluates to 9.
@@ 3.4 Numerical Results
Our results are presented in Figures 2 and 3. The first figure displays our predictions without considering the non-minimal interactions $(\kappa=0)$, while the second figure incorporates these interactions with $(\kappa=1)$. In both cases, several notable features of the Sommerfeld enhancement (SE) emerge that warrant discussion.
Firstly, as is well-known, SE is a velocity-dependent phenomenon. For instance, at a relative velocity of $(v=0.3)$ (dashed lines in Figures 2 and 3), which corresponds to the conditions of the early Universe when the relic density was established, no significant structure is observed. This is because the particles in this regime possess relativistically significant velocities, and the SE, being a non-relativistic effect, is not impacted under such conditions.
In contrast, for smaller velocities such as those expected within the galaxy $(v=10^{-3})$, pronounced oscillations and resonances become evident. We observed that the small oscillations depend on interactions mediated by the Higgs boson. Notably, these oscillations are attenuated for large values of the portal coupling constant $\lambda$. Interestingly, the inclusion of non-minimal interactions $(\kappa\neq 0$ introduces an additional resonance at a lower mass. However, the $\kappa$-terms do not significantly alter the magnitude of the oscillations or the amplitude of the resonances.
For masses that exhibit such resonances, approximately around $1.91$ TeV and $6$ TeV, the Sommerfeld factor experiences a substantial boost, increasing by nearly five orders of magnitude across all values of the portal coupling constant $\lambda$. This remarkable enhancement underscores the critical role of these parameters in shaping the indirect detection prospects for dark matter.
@@ 4 Photons Flux from Annihilation neutral Dark Matter particles
Using the expressions above, we calculated the photon flux resulting from the annihilation cross section (8) using the manual from [23], which serves as a method for the indirect detection of DM. This entails the integration of equation (29). Figures 4 and 5 depict the photon flux from VDM annihilation per second for $\kappa=0$ and $\kappa=1$ respectively. In all the cases we have used $\lambda=0.5$. We present the signals of photon flux resulting from the annihilation of VDM particles from the center of the galaxy for different final states ($\gamma\gamma$, $W^{+}W^{-}$, $ZZ$, and $hh$), with a solid angle of $1^{\circ}$, compared with the canonical cross-section with a mass of 100 GeV.Additionally, we show the total flux predicted at different values of the DM mass. We selected the masses to coincide with peaks and local minimums in the SE factor. We compare our results with Fermi-LAT sesitivity and the projected CTA sensitivity [20].
Figure 4: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, considering a region of the center of the Milky Way with solid angle $\Delta\Omega=0.96\times 10^{-3}$ sr (see Table 2 in [24]). This photon flux is measured at the center of the galaxy. This photon flux is measured at the center of the galaxy, with $\kappa=0$. We show differential sensitivity of CTA [20] and Fermi-Lat [25] in the galactic center.
Figure 5: Photons Flux from Annihilation of neutral Dark Matter particles. Where it is done for the Navarro-Frenk-White (NFW) profile, considering a region of the center of the Milky Way with solid angle $\Delta\Omega=0.96\times 10^{-3}$ sr (see Table 2 in [24]). This photon flux is measured at the center of the galaxy, with $\kappa=1$. We show differential sensitivity of CTA [20] and Fermi-Lat [25] in the galactic center.
@@ 5 Summary
In this paper, we have studied the Sommerfeld enhancement in a model where the DM candidate is a neutral component of a new massive vector field in the fundamental representation of $SU(2)_{L}$. We found that the annihilation cross-section predicted by our model is generally bigger than the one predicted in the cases of a spin $1/2$ doublet and a spin $1$ triplet. Additionally, the presence of non-minimal interactions of the doublet with the standard gauge model produces a richer resonance structure. Our results show that certain regions of the parameter space of our mode are testable by facilities like CTA.